# Multiple stellar populations in MOCCA globular cluster models: Transient spatial over-concentration of pristine red giant stars driven by strong dynamical encounters

### v1

@@ 4.2 Preferential ejection of enriched RGB progenitors by strong dynamical encounters
To explain the spatial over-concentration of $\rm 1P$ RGB stars at 13 Gyr, we first identified stars that evolve into RGBs at this epoch. These stars are near the turn-off mass, which is 0.8154 $\rm M_{\odot}$ for $\rm Z=0.001$ and had zero-age MS mass of approximately 0.83 $\rm M_{\odot}$. Assuming that all stars in the initial model (or at $\sim$100 Myr, when $\rm 2P$ is fully added) with masses between 0.81 and 0.83 $\rm M_{\odot}$ will have evolved into RGB stars by 13 Gyr, we define this mass range as the present-day RGB progenitors and examined those that escaped before 13 Gyr.
For $\rm 1P$ stars, we counted all escaping RGB progenitors within the 0.81–0.83 $\rm M_{\odot}$ mass range, regardless of the cause of escape, and found that 5048 $\rm 1P$ RGB progenitors had escaped the cluster. Similarly, for $\rm 2P$ stars, 1167 RGB progenitors had escaped. This significant difference is expected, as $\rm 1P$ was initially tidally filling. However, when considering only escapers due to strong dynamical interactions (binary-single or binary-binary encounters), the numbers shift dramatically: $\rm 1P$ RGB progenitor escapers drop to 150, while $\rm 2P$ increases to 403. Thus, $\rm 2P$ RGB progenitors are 2.7 times more likely to be dynamically ejected than $\rm 1P$ progenitors, consistent with their initial central concentration.
We find that $\rm 2P$ RGB progenitors are more frequently ejected from the inner cluster regions due to strong dynamical encounters. Matching escaping stars with interaction data shows that most present-day $\rm 2P$ RGB progenitors were ejected between 2 and 6 Gyr, primarily from within a few times the half-light radius ($R_{hl}$). After 6 Gyr, following the cluster’s migration to a larger Galactocentric radius, the ejection rate decreases, though $\rm 2P$ RGB progenitors continue to be preferentially removed. This trend is illustrated in Fig. 3, where the lower panel shows that most $\rm 2P$ escapers originate from the inner regions, especially before 6 Gyr, while the top panel histogram highlights the higher escape rate of $\rm 2P$ RGB progenitors compared to $\rm 1P$. These results confirm that strong dynamical encounters selectively remove $\rm 2P$ RGB progenitors, contributing to the observed over-concentration of $\rm 1P$ RGB stars at 13 Gyr.
To quantify this effect, we define the ratio:
@@
where $N_{\text{RGB,in\_cluster}}$ is the number of RGB stars present in the cluster at a given snapshot, and $N_{\text{RGB,progenitor\_dyn\_escapers}}$ is the cumulative number of RGB progenitors lost to strong encounters by that time. For $\rm 2P$ at 13 Gyr, we find $R\approx 1.06$ (430/402), meaning that the number of remaining RGB stars is nearly equal to those lost through strong encounters. At earlier times, such as 12.5 Gyr, this ratio is higher ($R\approx 1.13$), as fewer progenitors have escaped relative to in-cluster RGB stars. By 13.5 Gyr, the ratio returns to $R\approx 1.13$, indicating that the over-concentration of $\rm 1P$ relative to $\rm 2P$ RGB stars is most pronounced at 13 Gyr, when $R$ is closest to unity. This is further amplified by the small-number statistics of RGB stars relative to the total stellar population. At 13 Gyr, the number of $\rm 2P$ RGB stars in the cluster (430) is nearly equal to the number of progenitors ejected due to strong encounters (402). This near-parity magnifies the impact of missing $\rm 2P$ RGB progenitors. At earlier times (12.5 Gyr), a larger fraction of $\rm 2P$ RGB stars remain in the cluster ($R>1.1$), mitigating this effect. However, at 13 Gyr, when $R\approx 1.06$, the influence of escaped progenitors dominates, leading to the over-concentration of $\rm 1P$ RGB stars. By 13.5 Gyr, the ratio increases again, erasing this over-concentration.
@@
4.3 Conclusion
Our study suggests that the spatial distribution and, potentially, the kinematic properties of $\rm MPOP$ may depend on the type of stars observed. This effect is particularly relevant for GCs with present-day masses of a few $10^{5}\ \rm M_{\odot}$, which have retained only 10–20% of their initial mass. Despite the significant mass loss, such clusters may appear dynamically young due to BHS heating and migration to larger Galactocentric distances. When observations are limited to RGB stars, small-number statistics and dynamical interactions can distort their spatial distribution, leading to biased conclusions about the overall distribution of $\rm MPOP$. The over-concentration of $\rm 1P$ RGB stars relative to $\rm 2P$ RGB stars appears to be a transient effect.
Confirming the findings of Leitinger et al. (2023) and Cadelano et al. (2024), that in some MW GCs, $\rm 1P$ RGB stars are more centrally concentrated than $\rm 2P$ RGB stars, requires observations of MS stars. Our mocca simulations show no $\rm 1P$ over-concentration relative to $\rm 2P$ among MS stars. Verifying this observationally would be a crucial step toward understanding $\rm MPOP$ properties and supporting the AGB scenario of $\rm MPOP$ formation. Additionally, our results align with the kinematic analysis by Cadelano et al. (2024), which suggests that $\rm 2P$ stars initially formed more centrally concentrated than $\rm 1P$ stars, influencing their present-day spatial and kinematic distributions.
We have shown that the observed over-concentration of $\rm 1P$ relative to $\rm 2P$ can be explained within the AGB scenario, provided certain internal and external conditions are met. This effect appears to be transient, driven by fluctuations in the number and spatial distribution of RGB stars. Further analysis is needed to confirm these results, particularly regarding binary fraction distributions and velocity anisotropy in $\rm 1P$ and $\rm 2P$. In an upcoming study, we will conduct a more detailed comparison of mocca model results with observations from Leitinger et al. (2024) and Cadelano et al. (2024) to further investigate these trends.
Acknowledgments
MG, AH, GW, LH were supported by the Polish National Science Center (NCN) through the grant 2021/41/B/ST9/01191. AA acknowledges support for this paper from project No. 2021/43/P/ST9/03167 co-funded by the Polish National Science Center (NCN) and the European Union Framework Programme for Research and Innovation Horizon 2020 under the Marie Skłodowska-Curie grant agreement No. 945339. For the purpose of Open Access, the authors have applied for a CC-BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission.
@@ Acknowledgments
The data for the mocca model described in this paper, including the 13 Gyr snapshot, is available at https://doi.org/10.5281/zenodo.14908094. Additional data can be provided upon request to the corresponding author.

### v2

@@ 4.2 Preferential ejection of enriched RGB progenitors by strong dynamical encounters
To explain the spatial overconcentration of $\rm 1P$ RGB stars at 13 Gyr, we first identified stars that evolve into RGBs at this epoch. These stars are near the turn-off mass, which is 0.8154 $\rm M_{\odot}$ for $\rm Z=0.001$ and had a zero-age MS mass of approximately 0.83 $\rm M_{\odot}$. Assuming that all stars in the initial model (or at $\sim$100 Myr, when $\rm 2P$ is fully added) with masses between 0.81 and 0.83 $\rm M_{\odot}$ will have evolved into RGB stars by 13 Gyr, we defined this mass range as the present-day RGB progenitors and examined those that escaped before 13 Gyr.
For $\rm 1P$ stars, we counted all escaping RGB progenitors within the 0.81–0.83 $\rm M_{\odot}$ mass range, regardless of the cause of escape, and found that 5048 $\rm 1P$ RGB progenitors had escaped the cluster. Similarly, for $\rm 2P$ stars, 1167 RGB progenitors had escaped. This significant difference is expected, as $\rm 1P$ was initially tidally filling. However, when considering only escapers due to strong dynamical interactions (binary-single or binary-binary encounters), the numbers shift dramatically: $\rm 1P$ RGB progenitor escapers drop to 159, while $\rm 2P$ increases to 414. Thus, $\rm 2P$ RGB progenitors are 2.6 times more likely to be dynamically ejected than $\rm 1P$ progenitors, which is consistent with their initial central concentration.
We find that $\rm 2P$ RGB progenitors are more frequently ejected from the inner regions of the cluster via strong dynamical encounters. As is shown in Fig. 3, most $\rm 2P$ escapers are removed between 2 and 6 Gyr, primarily from within a few times the half-light radius ($R_{hl}$). After 6 Gyr, following the cluster’s migration to a larger Galactocentric distance, the ejection rate declines but remains higher for $\rm 2P$ stars. The lower panel shows that 2P escapers are mostly launched from the inner regions, while the top histogram highlights their higher escape rate compared to $\rm 1P$. Markers in the figure also reveal that a significant fraction of these ejections – especially for 2P stars – involve interactions with BHs (23.2% for 2P vs. 12.6% for 1P). Together, these results confirm that strong dynamical encounters, often involving BHs, preferentially remove $\rm 2P$ RGB progenitors, contributing to the overconcentration of $\rm 1P$ RGB stars at 13 Gyr.
To quantify this effect, we defined the ratio:
@@
where $N_{\text{RGB,in\_cluster}}$ is the number of RGB stars present in the cluster at a given snapshot, and $N_{\text{RGB,progenitor\_dyn\_escapers}}$ is the cumulative number of RGB progenitors lost to strong encounters by that time. For $\rm 2P$ at 13 Gyr, we find $R=1.04$ (430/414), indicating that the number of remaining RGB stars is nearly equal to those lost through strong encounters. At earlier times, such as 12.5 Gyr, this ratio is higher ($R=1.10$), as fewer progenitors have escaped relative to those still in the cluster. By 13.5 Gyr, the ratio rises to $R=1.11$, showing that the overconcentration of $\rm 1P$ relative to $\rm 2P$ RGB stars is most pronounced at 13 Gyr, when $R$ is closest to unity. This near-parity magnifies the impact of missing $\rm 2P$ RGB progenitors on the observed spatial distributions. At 13 Gyr, the number of 2P RGB stars in the cluster (430) is nearly matched by the number of progenitors ejected via strong encounters (414). By contrast, at 12.5 Gyr, the higher ratio ($R=1.1$) means more 2P RGB stars remain in the cluster, lessening this effect. By 13.5 Gyr, $R$ increases again to 1.11, diminishing the overconcentration. This increase occurs because lower-mass RGB progenitors, more numerous due to the IMF and still present in the cluster, begin evolving onto the RGB at this stage, thereby increasing the numerator in the ratio.
@@
4.3 Conclusions
Our study suggests that the spatial distribution and, potentially, the kinematic properties of $\rm MPOP$ may depend on the type of stars observed. This effect is particularly relevant for GCs with present-day masses of a few $10^{5}\ \rm{M_{\odot}}$, which have retained only 10–20% of their initial mass. Despite the significant mass loss, such clusters may appear dynamically young due to BHS heating and migration to larger Galactocentric distances. When observations are limited to RGB stars, small-number statistics and dynamical interactions can distort their spatial distribution, leading to biased conclusions about the overall distribution of $\rm MPOP$. The overconcentration of $\rm 1P$ RGB stars relative to $\rm 2P$ RGB stars appears to be a transient effect.
Confirming the findings of Leitinger et al. (2023) and Cadelano et al. (2024), that in some MW GCs, $\rm 1P$ RGB stars are more centrally concentrated than $\rm 2P$ RGB stars, requires observations of MS stars. Our mocca simulations show no $\rm 1P$ overconcentration relative to $\rm 2P$ among MS stars. Verifying this observationally would be a crucial step toward understanding $\rm MPOP$ properties and supporting the AGB scenario, and possibly other scenarios, of $\rm MPOP$ formation. Additionally, our results align with the kinematic analysis by Cadelano et al. (2024), which suggests that $\rm 2P$ stars initially formed more centrally concentrated than $\rm 1P$ stars, influencing their present-day spatial and kinematic distributions.
We have shown that the observed overconcentration of $\rm 1P$ relative to $\rm 2P$ can be explained within the AGB scenario, provided certain internal and external conditions are met. This effect appears to be transient, driven by fluctuations in the number and spatial distribution of RGB stars. Further analysis is needed to confirm these results, particularly regarding binary fraction distributions and velocity anisotropy in $\rm 1P$ and $\rm 2P$. In an upcoming study, we shall conduct a more detailed comparison of mocca model results with observations from Leitinger et al. (2024) and Cadelano et al. (2024) to further investigate these trends.
Acknowledgements.
We are grateful to the anonymous reviewer for their insightful comments, which helped improve the manuscript. MG, AH, GW, LH were supported by the Polish National Science Center (NCN) through the grant 2021/41/B/ST9/01191. AA acknowledges support for this paper from project No. 2021/43/P/ST9/03167 co-funded by the Polish National Science Center (NCN) and the European Union Framework Programme for Research and Innovation Horizon 2020 under the Marie Skłodowska-Curie grant agreement No. 945339. AH acknowledges support by the IDUB grant 140/04/POB4/0006 at Adam Mickiewicz University in Poznan, Poland. For the purpose of Open Access, the authors have applied for a CC-BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission.
@@ Acknowledgements.
The data for the mocca model described in this paper, including the 13 Gyr snapshot, is available at https://doi.org/10.5281/zenodo.14908094. Additional data can be provided upon request to the corresponding authors.
Appendix A MOCCA code
This work is based on the numerical simulations performed with the mocca Monte Carlo code (Giersz 1998; Hypki & Giersz 2013; Giersz et al. 2013; Hypki et al. 2022, 2024; Giersz et al. 2024). mocca is an advanced code that performs full stellar and dynamical evolution of real-size star clusters up to the Hubble time. The implementation of stellar and binary evolution within the mocca code is based on the rapid population synthesis code bse code (Hurley et al. 2000, 2002) that has been strongly updated by Belloni et al. (2017, 2018), Banerjee et al. (2020), and Kamlah et al. (2022). Strong dynamical interactions in mocca are performed with fewbody code (Fregeau et al. 2004; Fregeau & Rasio 2007). mocca can follow the full dynamical and stellar evolution of $\rm MPOP$, allowing for delayed $\rm 2P$ formation relative to $\rm 1P$. In this scenario, $\rm 2P$ forms through the re-accretion of gas surrounding the newly formed cluster, which mixes with enriched material from the ejected envelopes of AGB stars. Additionally, mocca incorporates the effects of cluster migration, driven by the rapidly evolving gravitational potential of the early Galaxy and dynamical friction.
The stellar and binary evolution parameters used in the presented simulations are referred to as Model C in Kamlah et al. (2022). In short, the metallicity of both populations in all the simulated models was set to Z = 0.001. The updated treatment for the evolution of massive stars was used according to Tanikawa et al. (2020) together with improved treatment for mass loss due to stellar winds and the inclusion of pair and pulsational pair-instability supernova (Belczynski et al. 2016). The masses of BHs and NS were determined according to the rapid supernovae prescriptions from Fryer et al. (2012). NS natal kicks were sampled from a Maxwelllian distribution with $\sigma=265$ km/s (Hobbs et al. 2005). However, for BHs, these natal kicks were reduced according to the mass fallback prescription (Belczynski et al. 2002; Fryer et al. 2012). The formation of neutron stars with negligible natal kicks through electron-capture or accretion-induced supernova was also enabled. Another feature of these models is the inclusion of gravitational wave recoil kicks whenever two BHs merge (Baker et al. 2008; Morawski et al. 2018). Low birth spins were assumed for BHs, with values uniformly sampled between 0 and 0.1 (Fuller & Ma 2019). The orientation of the BH spin relative to the binary orbit was randomly distributed (Morawski et al. 2018). The full list of changes in the recent mocca code, together with their detailed description, can be found in Giersz et al. (2024).
Appendix B BH subsystem evolution
The BHS plays a critical role in the development of 1P overconcentration and evolves differently from that in single-population GCs that are initially tidally underfilling. Figure 4 shows the time evolution of the BHS-to-cluster mass ratio, highlighting five key phases.
Figure 4: Time evolution of the BH subsystem mass as a fraction of the total cluster mass. Labels 1–5 mark the distinct evolutionary phases described in Section B
Figure 5: Time evolution of Lagrangian radii (10%, 50%, and 70%) for BHs, showing the progressive central concentration of the BHS
In phase 1, the BHS mass fraction rises rapidly due to BH formation from massive 1P stars and concurrent 1P mass loss as the cluster expands and overfills its tidal radius. In phase 2, the mass fraction drops following 2P formation, which significantly increases the total mass while contributing few BHs. The evolution beyond this point is shaped by BHS mass segregation and contraction (see Fig. 5). In phase 3, this segregation fuels energy generation that drives enhanced stellar escape, causing the BHS mass fraction to rise again. The brief dip around 2 Gyr reflects the final stages of BHS formation and early BH ejections. Phase 4 marks a period of balanced evolution, with the cluster expanding and preferentially losing 1P stars. As the system is only mildly nTF, escape remains efficient, and the BHS mass fraction increases further. At 5.5 Gyr, the cluster migrates to a larger Galactocentric distance and becomes strongly nTF (phase 5), suppressing stellar escape and leading to gradual BHS depletion through internal interactions. While the present-day BHS mass fraction in our model (5–6%) is higher than values inferred for most Galactic GCs—which are typically below 1% (Weatherford et al. 2020; Dickson et al. 2024)—it is comparable to estimates for exceptional cases such as Omega Centauri (Zocchi et al. 2019; Baumgardt et al. 2019). For clusters like NGC 3201 and NGC 6101, the inferred BH mass fractions are typically lower, around 1–2%. However, we note that multi-mass modeling techniques used to derive these estimates can systematically underpredict BH mass fractions above  1%, particularly in dynamically young clusters (see Fig. 1 in Dickson et al. 2024). Although NGC 3201, similar to our model, has a large present-day half-mass radius and appears dynamically young, our simulation suggests that it underwent significant dynamical evolution prior to its migration to a larger Galactocentric distance.
Appendix C Observational core radius
The observational core radius, as calculated in MOCCA, can vary substantially between output times, particularly at late stages of cluster evolution. This variability arises from the small number of bright evolved stars that dominate the central surface brightness, making the derived core radius sensitive to their spatial distribution. Figure 6 shows the time evolution of the observational core radius in our model. While the long-term trend reflects the global dynamical state of the cluster, significant variations between output times are apparent –ranging from below 1 pc to over 6 pc. These fluctuations are not necessarily indicative of structural changes, but rather reflect changes in the central surface brightness profile caused by the positions of a few luminous stars. For example, at the output time shown in the main paper (13.002 Gyr), the observational core radius is 3.13 pc and the central surface brightness (CSB) is 610 L/pc2. In the immediately following output (13.006 Gyr), these values shift to 2.07 pc and 817 L/pc2, respectively—closer to the observational values reported for NGC 3201 in the Harris (1996, updated 2010) catalog, which lists a core radius of 1.85 pc and a CSB of 914 L/pc2. Given these fluctuations, we emphasize that other structural parameters—such as total luminosity, cluster mass, and half-light radius—are more stable and provide a more reliable basis for comparison with observations. Additionally, the 3D core radius of the model at 13.002 Gyr is 1.22 pc, with a half-mass radius of 8.71 pc. At 13.006 Gyr, the 3D core radius is 1.77 pc, and the half-mass radius is 8.72 pc. These values are in excellent agreement with the parameters listed for NGC 3201 in the Baumgardt & Hilker (2018, updated 2024) catalog [^2] https://people.smp.uq.edu.au/HolgerBaumgardt/globular/parameter.html.
Figure 6: Observational core radius as a function of time.